OpenStudy (anonymous):

Find two positive real numbers whose product is a maximum. The sum is 10.

OpenStudy (asnaseer):

call then x and y, so you have: x + y = 10 y = 10 -x we can then write the product as: p = xy = x(10-x) = 10x - x^2 differentiate and set to zero to find what value of x gives the max. have you done differentials?